OpenStudy (anonymous):

what is the least common multiple of 3, 40 and 13?!?!

OpenStudy (anonymous):

c) 13 and 39. LCM = 39. GCD = 13. d) 5 and 8. LCM = 40. GCD = 1. e) 20 and 24. LCM = 120. GCD = 4 Finding the least common denominator for fractions is not as difficult as you might think. ... Note: In the examples below, we'll be adding three fractions instead of the usual two because the principles are the same. This will give you ... Multiples of 5 are 10, 15, 20, 25, 30, 35, 40,. ... 1/5 + 1/6 + 1/15 = 6/30 + 5/30 + 2/30 = 13/30 ...

OpenStudy (anonymous):

done

OpenStudy (anonymous):

can u help me with my question

OpenStudy (anonymous):

what is your question

OpenStudy (anonymous):

How does your body stay cool?

OpenStudy (anonymous):

do u know

OpenStudy (anonymous):

umm heat escapes from your head... Maybe haha I dont know!